OpenStudy (anonymous):

Pleases solve and explain : 2^(5/2) - 2^(3/2) =

OpenStudy (anonymous):

Ok, First tell me.. What is \[2^\frac{4}{2} = ?\]

OpenStudy (anonymous):

I forget the rules of exponents.

OpenStudy (anonymous):

Well, what is \[\frac{4}{2} = ?\]

OpenStudy (anonymous):

I forget the rules of exponents.

OpenStudy (anonymous):

2

OpenStudy (anonymous):

Right. So \[2^\frac{4}{2} = 2^2 = ?\]

OpenStudy (anonymous):

I forget the rules of exponents.

OpenStudy (anonymous):

But the exponents are 5/2 and 3/2. What does 4/2 have to do with the equation?

OpenStudy (anonymous):

Well that's a good question.. But we know that \(\frac{5}{2} = \frac{4}{2} + \frac{1}{2}\) And \(\frac{3}{2} = \frac{2}{2}+ \frac{1}{2}\) Right? And we know that \[a^{b+c} = a^b\cdot a^c\] Can we put these things together somehow?

OpenStudy (anonymous):

I forget the rules of exponents.

OpenStudy (anonymous):

5/2 also = 3/2+2/2 . And I've lost you on the following equation. But I believe that is one of the rules. Please continue.

OpenStudy (anonymous):

The last equation is a very important rule of exponents. When you multiply two powers of the same base you add their exponents. We can also do this in reverse.

OpenStudy (anonymous):

So: \[2^\frac{5}{2} - 2^\frac{3}{2}\]\[=2^\frac{4}{2}\cdot 2^\frac{1}{2} - 2^\frac{2}{2}\cdot 2^\frac{1}{2} \]

myininaya (myininaya):

\[2^{\frac{4}{2}+\frac{1}{2}}-2^{\frac{2}{2}+\frac{1}{2}}\] so this is your equation see if you can simplify this

myininaya (myininaya):

lol go polpak!

OpenStudy (anonymous):

Sorry yes, Myinin's expression is the step you should have in between my two.

myininaya (myininaya):

omg i said equation lol

myininaya (myininaya):

yes it is an expression

myininaya (myininaya):

rae what is the very next step after polpak's expression?